# Local-Cloud Inference Offloading for LLMs in Multi-Modal, Multi-Task, Multi-Dialogue Settings

### v1

@@ IV-C Main Experimental Results
Table I presents the main results of this study, with key takeaways described from four perspectives.
@@
Comparison with Naive Baselines. Compared to the local baseline, the proposed methods can significantly improve the response score by sacrificing latency and usage costs, thus enhancing the overall reward. Compared to the cloud baseline, we also boost the response score and conserve resources by selecting the appropriate modalities, which leads to improved rewards. In contrast to Random, our approach strategically explores the relationships between tasks/dialogues and LLMs/modalities, thereby optimizing decision-making processes to better align with the specific requirements and constraints of each task.
@@
Comparison with LLM-as-Agent Baselines Inspired by recent advances in LLM-as-agent [56, 57], we replace the RL component in our LCIO system with an LLM for decision making. Specifically, we design a system prompt to specify the background information, including the objective, output format, and possible actions. The workflow then follows an RL-like procedure: when faced with a particular state, we first convert it into a natural language prompt, after which the LLM agent outputs an action. Our goal is to evaluate the zero-shot LLM agent’s ability to consider modality-task relationships, understand context, and handle numerically sensitive tasks (e.g., avoiding resource constraint violations). To ensure a fair comparison, we do not factor in the latency or cost of the LLM agent here. As shown in Table I, different LLM agents exhibit substantial variability in performance. Some agents ignore all modalities (e.g., Mistral-7B-v0.3, Gemma-2-2b, Gemma-2-9b, GPT-4o-mini, GPT-4o, OpenAI o1-mini, OpenAI o1, and OpenAI o3-mini), while others choose to upload all available modalities for each task (e.g., FLAN-T5-large). Although some agents (e.g., Phi-3-mini, Phi-3.5-mini, and FLAN-T5-xl) appear to select different LLMs and modalities without violating any constraints, their response scores and overall rewards do not confer significant advantages. Thus, these zero-shot LLM agents do not demonstrate clear benefits. One reason is that they are not fine-tuned in the training dataset. Another reason is that, as generative models, LLM agents lack precise numerical calculation capabilities. Consequently, some agents violate constraints, whereas others adopt overly conservative strategies.
@@
Comparison with SOTA Baselines. Existing works do not address multi-task and multi-dialogue settings in a multi-modal scenario. For example, [33] utilizes the HumanEval dataset [58], which focuses on generating Python code based on task descriptions. This dataset uses a fixed evaluation metric, Pass@k, which measures the success rate of passing predefined unit tests. In contrast, our method considers a multi-modal, multi-task, and multi-dialogue setting, introducing challenges related to response score evaluation and uncertainty. To adapt methods such as AIwRG [33], which employs active inference, and PerLLM [49], which uses an upper confidence bound (UCB)-based approach, to our setting, we incorporate certain designs from this paper, including the unique mechanisms for LLM and modality selection. In addition, we replace the Pass@k metric with the average response score for each action in the dataset as the success metric. The results demonstrate that our LCIO system significantly outperforms these SOTA baselines. On the one hand, [33] fails to efficiently select LLMs and modalities, frequently opting to upload all three types of modality data sources for most actions. This results in substantial resource waste and redundancy in uploading unnecessary information to the cloud. On the other hand, [49] rarely considers any modality data, relying solely on the local LLM and text-only cloud LLM. Although this approach avoids violating any resource constraints, it leads to lower response scores, as the LLM lacks access to any information derived from modality data.
@@
Analysis of Our LCIO System. Most RL algorithms prefer to utilize the cloud LLM rather than the local LLM, as the gain in response score from the cloud LLM is much greater than the resource savings achieved by the local LLM. This indicates that GPT-4o, which is used for the cloud LLM, is much powerful than the Phi-3-mini used for the local LLM. Regarding modality selection, few RL algorithms choose a text-only cloud LLM, as it often results in interaction times comparable to or even longer than those with multi-modal inputs. This largely depends on the inference speed of the service provider rather than factors like network bandwidth. A potential reason is that text-only and multi-modal cloud LLMs may operate on different workflows and versions. However, in cases involving multi-modal inputs, we observe that single modality selection often receives preferential treatment by RL algorithms, as it provides valuable information with less latency and lower usage costs compared to selecting two or three modalities. RC-A2C demonstrates the best response scores and overall rewards while ensuring compliance with latency and usage cost constraints. It can be observed that RC-A2C selects as many modalities as possible to maximize the overall reward without violating resource constraints. Additionally, the synchronized updates and low-variance policy gradients of RC-A2C facilitate more stable and smooth policy updates. This method effectively adapts and optimizes resource constraints into the loss function, ensuring that these constraints are not violated.
IV-D Ablation Study
To further validate the effectiveness of our approach, in the last three rows of Table I, we utilize the top-performing RC-A2C as the backbone for ablation studies. Specifically, we remove the LLM selection, modality selection, and uncertain response score estimation components from RC-A2C, respectively. In the case of w/o LLM selection, an LLM is randomly selected, and if a cloud LLM is chosen, RC-A2C determines which modalities to use by sorting the action probabilities. For w/o modality selection, once RC-A2C selects the cloud LLM, the modalities are randomly chosen. For w/o response score estimation, we calculate the average response score of each action in the dataset for training, while still evaluating the test performance using our uncertain response score estimation method.
Compared to using RC-A2C, randomly selecting the LLM results in a degradation of the response score because, in many scenarios, the local LLM lacks the capability to complete the tasks. Randomly selecting the modalities leads to significant increases in latency and usage costs, as unnecessary modalities are uploaded, resulting in redundancy. Lastly, while using the average response score of actions may seem reasonable, it fails to account for the contextual information in the state. Specifically, the state includes historical modality upload information, which is crucial for response score estimation. When the response score estimation excludes this context, the estimation becomes biased, leading to incorrect guidance for the decision-making of the model, ultimately degrading its performance in task allocation and modality selection. The results demonstrate that LLM selection effectively manages tasks of varying difficulty by matching them with LLMs of appropriate capability, while modality selection enhances the alignment between modalities and tasks, thereby minimizing latency and usage costs by avoiding unnecessary uploads. Moreover, uncertain response score estimation leverages the contextual information to provide more accurate response score estimations, enabling the model to make well-informed decisions.
TABLE II: Trade-off between Metrics in Reward Function - Varying weights illustrate preferences for different metrics. More balanced weights results in a high response score and relatively low latency and usage cost.
@@ IV-D Ablation Study
Figure 6: Effect of Latency and Usage Cost Constraints. Left: Latency Constraint. Right: Usage Cost Constraint. LCIO demonstrates a faster reduction in resource constraint violations compared to baselines.
IV-E Trade-off Between Metrics
In Table II, we investigate the trade-offs among different metrics in the reward function by adjusting the weights for association, latency, and usage cost in Eq. (5), denoted as $\beta_{\Lambda}$, $\beta_{\psi}$, and $\beta_{\phi}$, respectively. We fix the weight of the response score to $\alpha=1$, as in Table I. We observe that neglecting any of these metrics leads to a degradation in the corresponding performance measure. For example, setting $\beta_{\Lambda}=0$ results in a lower response score because the association between the prompt and modalities is not considered. Furthermore, due to the presence of resource constraints, simply setting $\beta_{\psi}=0$ or $\beta_{\phi}=0$ not only causes high latency or high usage costs, respectively, but also does not provide a benefit to the response score. Although the RL model is penalized by the resource constraint in the loss function, it lacks explicit action-level penalties. As a result, the model may only account for cumulative costs, rather than recognizing the cost of individual actions. Concretely, uploading three modality data sources one at a time or uploading all three at once incurs roughly the same total resource cost. Consequently, when we balance the weights by setting $\beta_{\Lambda}=\frac{1}{3},\beta_{\psi}=\frac{1}{3},\beta_{\phi}=\frac{1}{3}$, we can attain a high response score while maintaining relatively lower latency and usage costs.
TABLE III: Local Devices Setting
@@ IV-E Trade-off Between Metrics
iPhone 15 (A16) 15.8 15 0.00096 6.69e-04
iPhone 15 Pro (A17 Pro) 35 15 0.00043 3.02e-04
This table details the inference time and usage costs required for local devices to perform a single inference using Phi-3-mini, as calculated by Eqs. (14) and (15). Note that in practical applications, these values are higher due to factors such as other task demands, memory bandwidth, I/O operations, utilization rates, power constraints, etc.
@@
Figure 7: Effect of Local Device Performance. Left and Right show the same experiment, demonstrating four different metrics. LCIO achieves superior performance on all local device compared to baselines.
IV-F Effect of Resource Constraints
In Fig. 6, we evaluate the performance of different schemes under varying latency and usage cost constraints, reflecting real-world scenarios where applications operate under different budgetary considerations. It is observed that as the settings for latency and usage cost constraints increase, all methods show a reduction in violations of these constraints, with RC-A2C achieving a faster reduction. This improvement is due to the resource constraints that are incorporated in the loss function to balance long-term cumulative rewards. In all different constraint settings, the proposed approach achieves the highest overall rewards, demonstrating its effectiveness in practical system deployments by considering response quality, latency and usage cost altogether. It is noted that even under extremely resource constraints, RC-A2C struggles to balance the trade-off between response score and resource costs effectively. Thus, in practical system deployments, besides employing the resource constraints to optimize long-term cumulative gains, it is also essential to establish safety layers (such as predicting the interaction time and usage cost and interventions directing RL to select local LLM) to mandatory avoid unforeseen high latency and usage costs, especially in time-sensitive applications like autonomous driving.
IV-G Effect of Different Local Device
To further confirm the advantage of our approach, in Fig. 7, we explore the effects of using different types of local devices, including Raspberry Pi, NVIDIA Jetson, and Apple iPhone. Detailed specifications of devices are available in Table III. We infer that the performance of different devices does not significantly affect the operation of the LCIO system, since their response scores and overall rewards do not show statistical significance. The underlying rationality is that for any device, latency and usage cost are normalized through min-max scaling and incorporated as part of the overall reward. The results are consistent with those in Table I, showing that our approach achieves (i) significant latency and cost savings compared to the cloud LLM baseline and (ii) significant improvements in response quality compared to the local LLM baseline, leading to the highest overall reward. Therefore, when the latency and usage costs of a local LLM are substantially lower than those of a cloud LLM, the LCIO system can operate effectively on any device. This precondition forms part of the background for the LCIO system.
Figure 8: Effect of Response Quality of Local LLM. Left: Response Score. Right: Overall Reward. LCIO consistently outperforms baselines across various deployments of local and cloud LLMs.
IV-H Effect of Local LLM Response Quality
@@
Finally, we evaluate the performance of the LCIO system by simulating various response qualities of the local LLM. Fig. 8 illustrates the results, where we adjust the response scores of local LLMs in the M4A1 dataset by adding or subtracting specific values (i.e., gap in Fig. 8) to simulate different local LLM capabilities. It is observed that regardless of the response quality of both the local LLM and the cloud LLM, the LCIO system consistently outperforms the three baseline models. This shows that LCIO can operate effectively in any combination of local and cloud LLMs. When the quality of the local LLM response is significantly lower than that of the cloud LLM, LCIO invariably opts for the cloud LLM and proceeds with modality selection. Conversely, as the response quality of the local LLM approaches that of the cloud LLM, LCIO begins to balance between the local and cloud options. An extreme scenario would be using the same LLM both locally and in the cloud, LCIO still capable of delivering superior response quality. This holds under the assumption that the local LLM does not support multi-modal inputs (or, considering a broader perspective, the cloud LLM cannot access user multi-modal data due to privacy concerns), LCIO can leverage the specific needs of certain tasks that require multi-modal data sources for better understanding of the environment. The overall results confirm the applicability of our methodology across a range of local LLM qualities.
V Conclusion
In this paper, we developed LCIO, the first local-cloud LLM inference offloading system designed for multi-modal, multi-task, and multi-dialogue scenarios. We employed RCRL to design the LLM selection and modality selection process to maximize long-term cumulative reward under resource constraints. Furthermore, we constructed a new dataset, M4A1, to evaluate the performance of the LCIO system under various conditions, including different RLs, reward trade-offs, resource constraints, devices, and network scenarios.
Future work will explore collaborative inference offloading among multiple users by developing frameworks that enable resource sharing, cooperative decision-making, and task partitioning to enhance system-wide efficiency and scalability. Additionally, we will investigate further options when a user is faced with multiple cloud service providers, focusing on adaptive strategies that dynamically select the optimal provider based on factors such as latency, usage cost, data privacy, and service reliability in real-time scenarios.

### v3

@@ 4.3. Main Experimental Results
Table 1 presents the main results of this study, with key takeaways described from four perspectives.
@@
Comparison with Naive Baselines. Compared to the local baseline, the proposed methods can significantly improve the response score by sacrificing latency and usage cost, thus enhancing the overall reward. Compared to the cloud baseline, we also boost the response score and conserve resources by selecting the appropriate modalities, which leads to improved rewards. In contrast to the random baseline, our approach strategically explores the relationships between tasks/dialogues and LLMs/modalities, thereby optimizing decision-making processes to better align with the specific requirements and constraints of each task.
@@
Comparison with SOTA LLM-as-Agent Baselines. We replace the RL agent in TMO with an LLM agent for decision-making processes. Our goal is to evaluate the zero-shot LLM agent’s ability to consider task-modality associations, understand contexts, and handle numerically sensitive tasks (e.g., avoiding resource constraint violations). As shown in Table 1, different LLM agents exhibit substantial variability in performance. Some agents ignore all modalities (e.g., Mistral-7B-v0.3, Gemma-2-2b, Gemma-2-9b, GPT-4o-mini, GPT-4o, OpenAI o1-mini, OpenAI o1, and OpenAI o3-mini), while others choose to upload all available modalities for each task (e.g., FLAN-T5-large). Although some agents (e.g., Phi-3-mini, Phi-3.5-mini, and FLAN-T5-xl) appear to select different LLMs and modalities without violating any constraints, their response scores and overall rewards do not confer significant advantages. Thus, these zero-shot LLM agents do not demonstrate clear benefits. One reason is that they are not fine-tuned in the training dataset. Another reason is that, as generative models, LLM agents lack precise numerical calculation capabilities. Consequently, some agents violate constraints, whereas others adopt overly conservative strategies.
@@
Comparison with SOTA Exploration-Decision Baselines. Existing works do not consider multi-task and multi-dialogue settings in a multi-modal scenario, and cannot be trivially deployed to our setting. For example, (He et al., 2024) utilize a benchmark with a fixed evaluation metric Pass@k that measures success rates. In contrast, our method addresses a multi-modal multi-task and multi-dialogue setting which introduces significant challenges related to response score evaluation and uncertainty. From the results, our TMO system significantly outperforms these SOTA baselines. (He et al., 2024) fails to efficiently select LLMs and modalities, frequently opting to upload all three types of modality data sources for most actions. This results in substantial resource waste and redundancy in uploading unnecessary information to the cloud. Conversely, (Yang et al., 2024) rarely considers any modality data, relying solely on the local LLM and text-only cloud LLM. Although this approach avoids violating any resource constraints, it leads to lower response scores, as the LLM lacks access to any information derived from modality data.
@@
Analysis of Our TMO System. Most RL algorithms prefer to utilize the cloud LLM rather than the local LLM, as the gain in response score from the cloud LLM is much greater than the resource savings achieved by the local LLM. Regarding modality selection, few RL algorithms choose a text-only cloud LLM, as it often results in interaction times comparable to or even longer than those with multi-modal inputs. This largely depends on the inference speed of the service provider rather than factors like network bandwidth. A potential reason is that text-only and multi-modal cloud LLMs may operate on different workflows and versions. However, in cases involving multi-modal inputs, we observe that single modality selection often receives preferential treatment by RL algorithms, as it provides valuable information with less latency and lower usage costs compared to selecting two or three modalities. RC-A2C demonstrates the best response scores and overall rewards while ensuring compliance with latency and usage cost constraints.
4.4. Ablation Studies
To validate our approach, we conduct ablation studies (Table 1, last three rows) by removing key components: RQ1 – LLM selection (substituting with random LLM selection), RQ2 – modality selection (using random modality selection instead), and RQ3 – uncertain response score estimation (using average action scores). Random LLM selection degrades response scores as local LLMs often lack task capabilities. Random LLM selection degrades response scores as local LLMs often lack task capabilities. Lastly, using average response scores fails to account for contextual information, particularly historical modality uploads, leading to biased estimations and degraded performance. Our results demonstrate that effective LLM selection matches tasks with appropriately capable LLMs, while modality selection enhances modality-task alignment, minimizing latency and costs. Additionally, uncertain response score estimation leverages contextual information for more accurate estimations and informed decisions.
Table 2. Trade-off between Metrics in Reward Function.
@@ 4.4. Ablation Studies
(a) Latency Constraint
(b) Usage Cost Constraint
Figure 4. Effect of Constraints. TMO demonstrates a faster reduction in resource constraint violations compared to baselines.
4.5. Trade-off Between Metrics
In Table 2, we investigate the trade-offs among different metrics in the reward function by adjusting the weights for association, latency, and usage cost in Eq. (5), denoted as $\beta_{\Lambda}$, $\beta_{\psi}$, and $\beta_{\phi}$, respectively. We fix the weight of the response score to $\alpha=1$. Due to the presence of resource constraints, simply setting $\beta_{\psi}=0$ or $\beta_{\phi}=0$ not only causes high latency or high usage cost, respectively, but also does not provide a benefit to the response score. Although the RL policy is penalized by the resource constraint in the loss function, it lacks explicit action-level penalties. As a result, the policy may only account for cumulative costs rather than recognizing individual action expenses, because uploading three modality data sources sequentially or simultaneously incurs roughly the same total resource cost. Consequently, when we balance the weights by setting $\beta_{\Lambda}=\frac{1}{3},\beta_{\psi}=\frac{1}{3},\beta_{\phi}=\frac{1}{3}$, we can attain a high response score while maintaining relatively lower latency and usage costs.
4.6. Effect of Resource Constraints
In Fig. 4, we evaluate the performance of different methods under varying latency and usage cost constraints, reflecting real-world scenarios where applications operate under different budgetary considerations. As latency and usage cost constraints increase, all methods demonstrate reduced constraint violations, with ours achieving more rapid reduction. This improvement stems from resource constraints incorporated into the loss function to balance the long-term cumulative rewards. In all different constraint settings, the proposed approach achieves the highest overall rewards, demonstrating its effectiveness in practical system deployments by considering response quality, latency, and usage cost altogether. However, under extremely limited resources, ours still struggles to effectively balance the trade-off between response score and resource costs. Therefore, in practical system deployments, beyond employing resource constraints to optimize long-term cumulative reward, establishing safety layers is also essential.
4.7. Effect of Different Local Devices
To further confirm the advantage of our approach, in Fig. 5, we explore the effects of using different types of local devices, with detailed specifications available in Table 3. We infer that the performance of different devices does not significantly affect the operation of the TMO system, since their response scores and overall rewards do not show statistical significance. The underlying rationality is that for any device, latency and usage cost are normalized through min-max scaling and incorporated as part of the overall reward. The results are consistent with those in Table 1, showing that our approach achieves (i) significant latency and cost savings compared to the cloud LLM baseline and (ii) significant improvements in response quality compared to the local LLM baseline, leading to the highest overall reward. Therefore, when the latency and usage cost of a local LLM are substantially lower than those of a cloud LLM, the TMO system can operate effectively on any device, a precondition that forms part of the background for the TMO system.
Table 3. Local Devices Setting.
@@ 4.7. Effect of Different Local Devices
iPhone 15 (A16) 15.8 15 0.00096 6.69e-7
iPhone 15 Pro (A17 Pro) 35 15 0.00043 3.02e-7
(a) Response Score & Overall Reward
(b) Latency & Usage Cost
@@
Figure 5. Effect of local device performance. TMO achieves superior performance on all local device compared to baselines.
4.8. Effect of Local LLM Response Quality
@@
Finally, we evaluate the TMO system’s performance by simulating varying local LLM response qualities. Fig. 6 demonstrates these results, wherein we modify the response scores in the M4A1 dataset by applying specific adjustments (represented as the “gap” in the figure) to simulate different local LLM capabilities. The TMO system consistently outperforms all three baselines regardless of local and cloud LLM response quality, demonstrating its effectiveness across any combination of local and cloud LLMs. When the local LLM response quality is substantially lower than the cloud LLM quality, TMO consistently selects the cloud LLM and proceeds with the modality selection. As the quality of the local LLM response approaches the quality of the cloud LLM, TMO balances between the local and cloud options. An extreme scenario would be using identical LLMs locally and in the cloud, TMO is still capable of delivering a superior response quality. This holds under the assumption that the local LLM does not support multi-modal inputs (or, considering a broader perspective, the cloud LLM cannot access user multi-modal data due to privacy concerns), which reflects a distinctive challenge described in RQ1. These results confirm our methodology’s applicability across diverse local LLM levels.
(a) Response Score
(b) Overall Reward
Figure 6. Effect of response quality of local LLM. TMO outperforms baselines across various deployments of local and cloud LLMs.
5. Conclusion
In this paper, we developed TMO, a novel local-cloud LLM inference offloading system designed specifically for multi-modal, multi-task, and multi-dialogue scenarios. We formulated the joint LLM and modality selection process as a resource-constrained reinforcement learning (RCRL) problem that maximizes long-term cumulative reward (response quality, latency, and usage cost) under practical user-defined constraints. Additionally, we curated a novel dataset, M4A1, enabling evaluation of TMO system performance across various configurations of modality availability, tasks and dialogues, resource constraints, and LLM response qualities. We demonstrated improvements in the overall reward obtained by TMO over several baselines. Future research will investigate multi-user collaborative inference through resource sharing and examine multi-server adaptive selection based on latency, cost, and other considerations.
Acknowledgements.
This work was supported in part by the National Science Foundation (NSF) under grant CPS-2313109, by the Office of Naval Research (ONR) under grant N00014-23-C-1016, and by the Air Force Office of Scientific Research (AFOSR) under grant FA9550-24-1-0083.